# A Data Perspective on Enhanced Identity Preservation for Diffusion Personalization

### v2

@@ 1 Introduction
We argue the reason that the previous methods fail to preserve identity details, e.g. text and logos, as shown in Figure 1, Figure 13, 14, 15 in Appendix 0.F, is because they cannot capture enough training information [12, 13, 32, 21, 9, 33, 30].
@@ 3 Method
Enhancing Training with Prompts. The inclusion of prompts in training instances plays a pivotal role in our approach. While employing a uniform prompt format such as “a [identifier_token] [class noun]” for all training examples may seem straightforward, it inadvertently increases the risk of associating the token with both foreground and background, causing overfitting to the training data. To address this challenge, we developed this idea further by incorporating background descriptions into each image, as depicted in the upper-right corner of Figure 2. Additionally, we replaced the provided class name with more specific ones, which will be discussed further in Appendix 0.B. It’s important to note that merely adding the background prompt alone does not yield the desired benefits. Instead, our contribution is on constructing a dedicated regularization dataset, which proves to be essential in addressing the issue of overfitting, as elaborated in subsequent sections.
Generating regularization examples against training prompts. We present a novel approach for generating a regularization dataset, which overcomes the ineffectiveness of the original example prompt. For instance, let’s consider the initial prompt “a backpack on a rock”. We enhance this prompt by introducing structural components in the format of “a [shape] [color] [texture] [class noun] [background]”, where [shape], [color], and [texture] represent randomly selected adjectives drawn from a pool of 100 options for each attribute. These adjectives are then combined to create prompts, such as “a contoured orchid woven backpack on a rock” and “a semicircular turquoise canvas backpack hanging on a tree branch”. The created prompts are used for image generation, as demonstrated in the lower-left corner of Figure 2. The pools of the adjectives are generated automatically by Large Language Models [19]. For more detailed information on this process, please refer to Appendix 0.A.
@@
Adaption for Living Entities. For living entities, we use descriptors like body, skin/fur, and emotion instead of shape, color, and texture. Additionally, we introduce motion into the backgrounds, transitioning from static scenes like “in an urban city” to dynamic contexts such as “walking in an urban city”. For a detailed explanation of this adaptation process, please refer to Appendix 0.A.
@@
Implementation Details We create 2000 regularization images for each subject, distributed as follows: 20% photorealistic images have different subjects in same background; 60% photorealistic images have different subjects in different background; and 20% differently styled images have different subjects in different background. We opt for the identifier word “olis” instead of the more commonly used “sks”. This choice is based on the fact that “olis” corresponds to the least frequently utilized token in the model’s vocabulary [1]. Each training batch contains one example from training set and one example from regularization set. For SD, we fine-tune the entire model with a learning rate of 2e-6 and perform inference using 200 steps of DDIM [31]. For SDXL, which has a larger model size, we employ a LoRA with a rank of 32 for both the text encoders and UNet. We also train the text embeddings. We set learning rate to 1e-4. We use 50 steps of DDIM for inference. Additional details regarding the optimal iteration range can be found in Appendix 0.E.
@@ 4.1 Result Comparison
The qualitative comparison is depicted in Figure 1 and Figure 3. We use “<new>” to represent the “ [identifier_token]”. Our approach outperforms alternative methods in preserving intricate subject-specific details, such as logos on drink cans, unique icons on backpacks, and fine hair textures on toy monsters.
Figure 3: Qualitative Results.Our results preserves finer details of the subject without overfitting to the training images or losing text alignment.
@@
In our quantitative assessment, we follow DreamBooth to employ DINO [4] and CLIP-I [22], to evaluate subject fidelity. For text alignment evaluation, we argue that the CLIP-T used in DreamBooth may yield inaccurate measurements of text-image alignment when vague class names are employed. As shown in Figure 4, there is a notable discrepancy between the utilization of vague class names, such as “toy”, and more specific object names, such as “duck toy”, on the ground truth images. Notably, the CLIP-T score appears to be significantly influenced by the nomenclature chosen for the object, and thereby potentially undermines its accuracy as an indicator of text-image alignment. To delve deeper into this matter, we calculate the CLIP-T score on the original images and the manually added prompts. Table 1 presents that, when using vague class names, the CLIP-T score for ground truth text-image pairs falls even below the conventional threshold of 0.3, which is typically considered as the threshold for assessing text-image pair compatibility [28]. To rectify this issue, we replace vague class names with highly specific object names, which results in a substantial improvement in the CLIP-T score for ground truth text-image pairs. Additional details regarding the modified nomenclature list can be found in Appendix 0.B.
We conducted a user study to compare with DreamBooth [25]. We asked 20 people to answer 30 questions about subject alignment and 30 questions about text alignment, in a total of 1200 questions. The questions are randomly sampled from a large pool of DreamBench test prompts. The details can be found in Appendix 0.C. As shown in Table 2, our method is overwhelmingly preferred for subject alignment and comparable in text alignment. We also compared the effect of a different base model with SDXL instead of SD. SDXL is more expressive in details and hence leads to better subject alignment. For textural alignment SD works slightly better.
@@ 4.2 Model Analysis
How does the training and regularization dataset size affects the performance? We conducted grid experiments on number of training examples 1, 2, 4, and number of regularization examples 100, 500, 1000, 2000. The results are illustrated in Figure  8. Our method is robust to smaller regularization set size, as only 100 regularization examples can also effectively prevent overfitting and the model is still able to preserve the very fine details of the subjects.
Is the random dropout of adjective words necessary? When altering attributes like color, we found that using random dropout is essential. Without it, the model tends to excessively preserve the original identity, and struggle in color changes. With random dropout, the task becomes easier while still preserving identity, as shown in Figure 9. It’s worth noting that for color modification, using “a <new>[color] [class noun]” works better than “a [color] <new>[class noun]”.
@@ 6 Conclusion
Appendix 0.A Formatted prompt generation
@@
  • background: give me 500 phrases that describe the background, such as “on the table", as diverse as possible.
@@
  • motion: give me 1000 different short concise sentences that contains a special token “\$concept" which stands for a specific animal, which can be a dog, a cat or a human. For example: “a \$concept sitting in a temple", “a \$concept walking in a supermarket". Keep “a \$concept" in the sentences.
After removing duplicated ones, there are 89 bodies, 86 skins/furs, 75 emotions, and 744 motions. For humans, we replace the word "animal" above with "person".
@@
  • style: give me 100 image style descriptions, such as “a photo of", and “a painting of".
@@
Appendix 0.B Better Category Naming
We show the name change in Table 3. As shown in Section 4.1, the CLIP-T score makes more sense in the changed name than the original name.
@@ Appendix 0.B Better Category Naming
Appendix 0.C Details of User Study
We randomly sampled and paired 300 comparisons of ours(SD) versus DreamBooth, half of which is for the subject alignment and the other half for the text alignment. For subject alignment, we randomly sampled a ground truth image and asked "The foreground object in which image is more similar to the reference?". For text alignment, we asked "Which image better depicts {}?", where {} is replaced by the prompt. We equally divided the questions into 10 groups. Each person randomly received one group. We did the same for ours(SDXL) versus ours(SD).
Appendix 0.D Extension: Using BLIP to Generate Captions
We tried to use BLIP [16] to generate more personalized captions for the training example images. BLIP outputs a caption for the input image and can be conditioned on the format. We condition BLIP so that it generates prompts that start with "a [subject]". For instance for the "tortoise plushie" image BLIP generates
@@ Appendix 0.D Extension: Using BLIP to Generate Captions
To unify the prompt format, we task ChatGPT to‘Change the following sentence to the format "A $<$new$>$ tortoise plushie blablabla". The "$<$new$>$" is a special token that needs to be inserted before tortoise plushie.’ The result are the following prompts
@@
The results of this "tortoise plushie" dataset is shown in Figure 14. With this addition of using BLIP, it alleviated writing the prompt examples manually, i.e., it replaced the manual steps in Section 3.
Appendix 0.E Optical Number of Iterations of DreamBench
We show the name change in Table 4. As shown in Section 4.1, the CLIP-T score makes more sense in the changed name than the original name.
@@ Appendix 0.E Optical Number of Iterations of DreamBench
Appendix 0.F More Results

### v4

@@ 1 Introduction
We argue the reason that the previous methods fail to preserve identity details, e.g. text and logos, as shown in Figure 1, Figure 14, 15, 16 in Appendix F, is because they cannot capture enough training information [12, 13, 32, 21, 9, 33, 30].
@@ 3 Method
Enhancing Training with Prompts. The inclusion of prompts in training instances plays a pivotal role in our approach. While employing a uniform prompt format such as “a [identifier_token] [class noun]” for all training examples may seem straightforward, it inadvertently increases the risk of associating the token with both foreground and background, causing overfitting to the training data. To address this challenge, we developed this idea further by incorporating background descriptions into each image, as depicted in the upper-right corner of Figure 2. Additionally, we replaced the provided class name with more specific ones, which will be discussed further in Appendix B. It’s important to note that merely adding the background prompt alone does not yield the desired benefits. Instead, our contribution is on constructing a dedicated regularization dataset, which proves to be essential in addressing the issue of overfitting, as elaborated in subsequent sections.
Generating regularization examples against training prompts. We present a novel approach for generating a regularization dataset, which overcomes the ineffectiveness of the original example prompt. For instance, let’s consider the initial prompt “a backpack on a rock”. We enhance this prompt by introducing structural components in the format of “a [shape] [color] [texture] [class noun] [background]”, where [shape], [color], and [texture] represent randomly selected adjectives drawn from a pool of 100 options for each attribute. These adjectives are then combined to create prompts, such as “a contoured orchid woven backpack on a rock” and “a semicircular turquoise canvas backpack hanging on a tree branch”. The created prompts are used for image generation, as demonstrated in the lower-left corner of Figure 2. The pools of the adjectives are generated automatically by Large Language Models [19]. For more detailed information on this process, please refer to Appendix A.
@@
Adaption for Living Entities. For living entities, we use descriptors like body, skin/fur, and emotion instead of shape, color, and texture. Additionally, we introduce motion into the backgrounds, transitioning from static scenes like “in an urban city” to dynamic contexts such as “walking in an urban city”. For a detailed explanation of this adaptation process, please refer to Appendix A.
@@
Implementation Details We create 2000 regularization images for each subject, distributed as follows: 20% photorealistic images have different subjects in same background; 60% photorealistic images have different subjects in different background; and 20% differently styled images have different subjects in different background. Additional details regarding the optimal iteration range can be found in Appendix E.
@@ 4.1 Result Comparison
The qualitative comparison is depicted in Figure 1 and Figure 10-16 in Appendix F. We use “new” to represent the “ [identifier_token]”. Our approach outperforms alternative methods in preserving intricate subject-specific details, such as logos on drink cans, unique icons on backpacks, and fine hair textures on toy monsters.
@@
In our quantitative assessment, we follow DreamBooth to employ DINO [4] and CLIP-I [22], to evaluate subject fidelity. For text alignment evaluation, we argue that the CLIP-T used in DreamBooth may yield inaccurate measurements of text-image alignment when vague class names are employed. More details and discussion can be found in Appendix B. To rectify this issue, we replace vague class names with highly specific object names, which results in a substantial improvement in the CLIP-T score for ground truth text-image pairs, as shown in Table 1.
We conducted a user study to compare with DreamBooth [25]. We asked 20 people to answer 30 questions about subject alignment and 30 questions about text alignment, in a total of 1200 questions. The questions are randomly sampled from a large pool of DreamBench test prompts. The details can be found in Appendix C. As shown in Table 2, our method is overwhelmingly preferred for subject alignment and comparable in text alignment. We also compared the effect of a different base model with SDXL instead of SD. SDXL is more expressive in details and hence leads to better subject alignment. For textural alignment SD works slightly better.
@@ 4.2 Model Analysis
How does the training and regularization dataset size affects the performance? We conducted grid experiments on number of training examples 1, 2, 4, and number of regularization examples 100, 500, 1000, 2000. The results are illustrated in Figure  8 in Appendix F. Our method is robust to smaller regularization set size, as only 100 regularization examples can also effectively prevent overfitting and the model is still able to preserve the very fine details of the subjects.
Is the random dropout of adjective words necessary? When altering attributes like color, we found that using random dropout is essential. Without it, the model tends to excessively preserve the original identity, and struggle in color changes. With random dropout, the task becomes easier while still preserving identity, as shown in Figure 9 in Appendix F. It’s worth noting that for color modification, using “a <new>[color] [class noun]” works better than “a [color] <new>[class noun]”.
@@ 6 Conclusion
Appendix A Formatted prompt generation
@@
  • background: give me 500 phrases that describe the background, such as “on the table”, as diverse as possible.
@@
  • motion: give me 1000 different short concise sentences that contains a special token “\$concept” which stands for a specific animal, which can be a dog, a cat or a human. For example: “a \$concept sitting in a temple”, “a \$concept walking in a supermarket”. Keep “a \$concept” in the sentences.
After removing duplicated ones, there are 89 bodies, 86 skins/furs, 75 emotions, and 744 motions. For humans, we replace the word ”animal” above with ”person”.
@@
  • style: give me 100 image style descriptions, such as “a photo of”, and “a painting of”.
@@
Appendix B Better Category Naming
We show the name change in Table 3. As shown in Figure 6, there is a notable discrepancy between the utilization of vague class names, such as “toy”, and more specific object names, such as “duck toy”, on the ground truth images. Notably, the CLIP-T score appears to be significantly influenced by the nomenclature chosen for the object, and thereby potentially undermines its accuracy as an indicator of text-image alignment. To delve deeper into this matter, we calculate the CLIP-T score on the original images and the manually added prompts. Table 1 presents that, when using vague class names, the CLIP-T score for ground truth text-image pairs falls even below the conventional threshold of 0.3, which is typically considered as the threshold for assessing text-image pair compatibility [28]. To rectify this issue, we replace vague class names with highly specific object names, which results in a substantial improvement in the CLIP-T score for ground truth text-image pairs. Additional details regarding the modified nomenclature list can be found in Appendix B.
@@ Appendix B Better Category Naming
Appendix C Details of User Study
We randomly sampled and paired 300 comparisons of ours(SD) versus DreamBooth, half of which is for the subject alignment and the other half for the text alignment. For subject alignment, we randomly sampled a ground truth image and asked ”The foreground object in which image is more similar to the reference?”. For text alignment, we asked ”Which image better depicts {}?”, where {} is replaced by the prompt. We equally divided the questions into 10 groups. Each person randomly received one group. We did the same for ours(SDXL) versus ours(SD). We provide an example of our interface in Figure 7.
Figure 7: An example of the user study interface. The left is an example for question of object alignment, and the right is an example for question of text alignment. Each user was asked to answer 30 questions about subject alignment (left) and 30 questions about text alignment (right). The examples in the questions are randomly sampled from a large pool.
Appendix D Extension: Using BLIP to Generate Captions
We tried to use BLIP [16] to generate more personalized captions for the training example images. BLIP outputs a caption for the input image and can be conditioned on the format. We condition BLIP so that it generates prompts that start with ”a [subject]”. For instance for the ”tortoise plushie” image BLIP generates
@@ Appendix D Extension: Using BLIP to Generate Captions
To unify the prompt format, we task ChatGPT to‘Change the following sentence to the format ”A $<$new$>$ tortoise plushie blablabla”. The ”$<$new$>$” is a special token that needs to be inserted before tortoise plushie.’ The result are the following prompts
@@
The results of this ”tortoise plushie” dataset is shown in Figure 15. With this addition of using BLIP, it alleviated writing the prompt examples manually, i.e., it replaced the manual steps in Section 3.
Appendix E Implementation Details
We opt for the identifier word “olis” instead of the more commonly used “sks”. This choice is based on the fact that “olis” corresponds to the least frequently utilized token in the model’s vocabulary [1]. Each training batch contains one example from training set and one example from regularization set. For SD, we fine-tune the entire model with a learning rate of 2e-6 and perform inference using 200 steps of DDIM [31]. For SDXL, which has a larger model size, we employ a LoRA with a rank of 32 for both the text encoders and UNet. We also train the text embeddings. We set learning rate to 1e-4. We use 50 steps of DDIM for inference. We show the best number of iterations in Table 4. For simplicity, we use 4000 and 8000 iterations for SD and SDXL, respectively.
@@ Appendix E Implementation Details
Appendix F More Results
Figure 10: More Results on inanimate objects.